OpenStudy (anonymous):

(1/4 a-b)(1/4a+b) multiply

OpenStudy (anonymous):

1/16a^2 - b^2

OpenStudy (anonymous):

how though???? khan hey ya

OpenStudy (saifoo.khan):

\[(\frac{a}{4}-b) (\frac{1}{4 a}+b)\]

OpenStudy (saifoo.khan):

Heyyy Rose! Supp?

OpenStudy (anonymous):

so whats my answer?

OpenStudy (anonymous):

you foil it because you have a plus and a minus you know that the middle terms cancel eachother out: (1/3a)(1/4a) + (1/4ab) - (1/4ab) + (b)(-b)

OpenStudy (anonymous):

the answer is the first thing i posted

OpenStudy (anonymous):

ok thank u

OpenStudy (saifoo.khan):

hmm.

OpenStudy (anonymous):

hmmm???

OpenStudy (anonymous):

hmm indeed. this is \[(\frac{1}{4}a-b)(\frac{1}{4}a+b)\] right?

OpenStudy (anonymous):

yes